# Category $\mathcal{O}$ for hybrid quantum groups and non-commutative Springer resolutions

### v3

@@ 2.3.4. Translation functors
For any $\omega_{1},\omega_{2}\in\Xi$, there is a unique dominant element $\nu$ in the $W$-orbit of $\omega_{2}-\omega_{1}$. Denote by $V(\nu)$ the Weyl module for $U_{\zeta}$ of highest weight $\nu$. We view $V(\nu)$ as a $U$-module via the natural map $U\rightarrow U_{\zeta}$. We define the translation functors by
$T_{\omega_{1}}^{\omega_{2}}:U\text{-}\mathrm{Mod}^{\Lambda,\omega_{1}}_{S}\rightarrow U\text{-}\mathrm{Mod}^{\Lambda,\omega_{2}}_{S},\quad M\mapsto\mathrm{pr}_{\omega_{2}}(M\otimes V(\nu)),$
$T^{\omega_{1}}_{\omega_{2}}:U\text{-}\mathrm{Mod}^{\Lambda,\omega_{2}}_{S}\rightarrow U\text{-}\mathrm{Mod}^{\Lambda,\omega_{1}}_{S},\quad M\mapsto\mathrm{pr}_{\omega_{1}}(M\otimes V(\nu)^{*}),$
where $\mathrm{pr}_{\omega_{i}}$ is the natural projection to the block $U\text{-}\mathrm{Mod}^{\Lambda,\omega_{i}}_{S}$. Note that $T_{\omega_{1}}^{\omega_{2}}$ and $T^{\omega_{1}}_{\omega_{2}}$ are exact and biadjoint to each other. The translation functors preserve the subcategory ${\mathcal{O}}^{U}_{S}$ and induce translation functors between blocks in ${\mathcal{O}}^{U}_{R}$, for any commutative Noetherian $S$-algebra $R$.
For any simple reflection $s\in{\mathrm{I}}_{\mathrm{af}}$, we fix an element $\omega_{s}\in\Xi_{\mathrm{sc}}$ whose stabilizer under the $W_{\mathrm{af}}$-action is $\{1,s\}$. We abbreviate $T^{0}_{s}=T^{0}_{\omega_{s}}$ and $T_{0}^{s}=T_{0}^{\omega_{s}}$. Define the reflection functor on $U\text{-}\mathrm{Mod}^{\Lambda,0}_{S}$ by
$\Theta^{r}_{s}:=T^{0}_{s}\circ T_{0}^{s}.$
[…]
@@
We similarly define translation functors and reflection functors on the categories ${\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{\chi_{\omega}}$ with $\omega\in\Xi$.
@@ 2.3.5. $\Lambda$-translations
Let $U^{\prime}$ be $U$ as in §2.3.3, or be $u_{\zeta}$, $u_{\zeta}U^{+}_{\zeta}$. Let $R$ be a deformation ring for $U^{\prime}$. For any $\nu\in\Lambda$, there is a trivial $U^{\prime}$-module ${\mathbb{C}}_{l\nu}$ supported on the weight $l\nu$. It gives auto-equivalences $-\otimes{\mathbb{C}}_{l\nu}$, $\nu\in\Lambda$ on $U^{\prime}\text{-}\mathrm{Mod}^{\Lambda}_{R}$ preserving each blocks, called the $\Lambda$-translations.
@@ 3.1. Non-commutative Springer resolutions
For any standard parabolic subgroup $P$ in $G$, we set $\tilde{{\mathfrak{g}}}_{P}=G\times^{P}\mathrm{Lie}(P)$ and abbreviate $\tilde{{\mathfrak{g}}}=\tilde{{\mathfrak{g}}}_{B}$. There is a natural map $\pi_{P}:\tilde{{\mathfrak{g}}}\rightarrow\tilde{{\mathfrak{g}}}_{P}$. The Grothendieck resolution is $\pi:\tilde{{\mathfrak{g}}}\rightarrow{\mathfrak{g}}$. If $P=P_{s}$ is the minimal parabolic subgroup associated to $s\in{\mathrm{I}}$, we will abbreviate $\tilde{{\mathfrak{g}}}_{s}=\tilde{{\mathfrak{g}}}_{P_{s}}$ and $\pi_{s}=\pi_{P_{s}}$. Denote by ${\mathcal{O}}_{\mathcal{B}}(\lambda)$ the line bundle $G\times^{B}{\mathbb{C}}_{\lambda}\rightarrow{\mathcal{B}}$, and denote by ${\mathcal{O}}_{Y}(\lambda)$ its pullback along any $Y\rightarrow{\mathcal{B}}$. Consider the ${\mathbb{C}}^{\times}$-action on ${\mathfrak{g}}$ via $t.x=t^{-2}x$, for any $t\in{\mathbb{C}}^{\times}$, $x\in{\mathfrak{g}}$. It induces ${\mathbb{C}}^{\times}$-actions on $\tilde{{\mathfrak{g}}}_{P}$ for any $P$.
@@ 3.1.1. Affine braid group actions
${{\mathcal{T}}or}_{i}^{{\mathcal{O}}_{\mathfrak{g}}}({\mathcal{O}}_{\tilde{{\mathfrak{g}}}},{\mathcal{O}}_{X})=0,\quad\forall i>0.$
We set $\tilde{{\mathfrak{g}}}_{P,X}:=\tilde{{\mathfrak{g}}}_{P}\times_{\mathfrak{g}}X$ and write $\pi_{P}=\pi_{P}\times_{\mathfrak{g}}X$ by abuse of notations. There is a functor $\Theta_{P}^{c}:=\pi_{P}^{*}\pi_{P*}$ on $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$. We abbreviate $\Theta^{c}_{s}=\Theta^{c}_{P_{s}}$ for any $s\in{\mathrm{I}}$, which are called the reflection functors.
@@
For any $s\in{\mathrm{I}}$, there is a self-equivalence ${\mathbb{T}}^{c}_{s}$ on $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$ with natural transformations
[…]
which is a distinguished triangle when applying on any object in $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$. The assignments
$\tilde{s}\mapsto({\mathbb{T}}^{c}_{s})^{-1}\quad\text{and}\quad\theta_{\lambda}\mapsto-\otimes^{L}_{{\mathcal{O}}_{\tilde{{\mathfrak{g}}}_{X}}}{\mathcal{O}}_{\tilde{{\mathfrak{g}}}_{X}}(\lambda),\qquad\forall s\in{\mathrm{I}},\ \forall\lambda\in\Lambda,$
define a (weak, right) $B_{\mathrm{ex}}$-action on $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$. Moreover, for any closed subgroup $K\subset G\times{\mathbb{C}}^{\times}$ and $K$-equivariant morphism $X\rightarrow{\mathfrak{g}}$, the constructions above are naturally $K$-equivariant.
@@ 3.1.2. Non-commutative Springer resolutions
Let $B^{+}_{\mathrm{ex}}$ be the semi-group of $B_{\mathrm{ex}}$ generated by the canonical lifts $\tilde{s}$ with $s\in{\mathrm{I}}_{\mathrm{af}}$.
@@
    There is a unique t-structure (called the exotic t-structure) on $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$ defined by
    $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})^{\leq}=\{{\mathcal{F}}|\ \pi_{*}(b{\mathcal{F}})\in D^{\leq}\mathrm{Coh}X,\ \forall b\in B^{+}_{\mathrm{ex}}\};$
    $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})^{\geq}=\{{\mathcal{F}}|\ \pi_{*}(b^{-1}{\mathcal{F}})\in D^{\geq}\mathrm{Coh}X,\ \forall b\in B^{+}_{\mathrm{ex}}\}.$
@@
    There exists a $G\times{\mathbb{C}}^{\times}$-equivariant tilting bundle ${\mathcal{E}}$ on $\tilde{{\mathfrak{g}}}$ such that
    $R\mathcal{H}om_{\tilde{{\mathfrak{g}}}_{X}}({\mathcal{E}}_{X},-):\ D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})\rightarrow D^{b}({\mathrm{A}}_{X}\text{-}\mathrm{mod})$
    is an equivalence, where ${\mathcal{E}}_{X}:={\mathcal{E}}\times_{\mathfrak{g}}X$, ${\mathrm{A}}_{X}={\mathrm{A}}\otimes_{{\mathcal{O}}_{\mathfrak{g}}}{\mathcal{O}}_{X}$ and ${\mathrm{A}}=\mathcal{E}nd_{\tilde{{\mathfrak{g}}}}^{\mathrm{op}}({\mathcal{E}})$, and such that this equivalence is t-exact with respect to the exotic t-structure on the LHS and the trivial t-structure on the RHS.
@@
    Moreover, for any closed subgroup $K\subset G\times{\mathbb{C}}^{\times}$ and $K$-equivariant morphism $X\rightarrow{\mathfrak{g}}$, we have similar statements for $D^{b}\mathrm{Coh}^{K}(\tilde{{\mathfrak{g}}}_{X})$.
By definition, the functor $\pi_{*}$ is t-exact for the exotic t-structure on $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$. It yields a functor $\pi_{*}:{\mathrm{A}}_{X}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}X$. By [12], the functors $\Theta^{c}_{P}$, $\Theta^{c}_{s}$ ($s\in{\mathrm{I}}_{\mathrm{af}}$) are also exact for the exotic t-structure. We thus view them as functors on ${\mathrm{A}}_{X}\text{-}\mathrm{mod}$.
@@
$D^{b}({\mathrm{A}}_{{\mathfrak{g}}_{\hat{0}}}\text{-}\mathrm{mod})\simeq D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{\hat{0}}),$
@@
(3.1) $D^{b}({\mathrm{A}}\text{-}\mathrm{mod}_{0})\simeq D^{b}\mathrm{Coh}_{{\mathcal{B}}}(\tilde{{\mathfrak{g}}}).$
@@ 3.2. Localization theorem for ${\mathfrak{U}}_{\zeta}$
The localization theorem for ${\mathfrak{U}}_{\zeta}$ was established by Backelin–Kremnizer [5] and Tanisaki [44], [45], [42]. In this subsection, we give a reminder on the localization theorem mainly following [42].
@@ 3.2.1. Quantum differential operators
Lunts—Rosenberg [28] introduced the quantized flag variety ${\mathcal{B}}_{\zeta}$ as a non-commutative scheme over ${\mathcal{B}}$. Let ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}$ be the pushforward of the structure sheave of ${\mathcal{B}}_{\zeta}$ on ${\mathcal{B}}$, which is a coherent sheaf of (in general non-commutative) algebras on ${\mathcal{B}}$. In [43], Tanisaki defined the ring of enhanced differential operators on ${\mathcal{B}}_{\zeta}$ (note that in loc. cit. the convention for flag variety is $B^{-}\backslash G$, which is different from us). Let $\widetilde{{\mathcal{D}}}_{\zeta}$ be its localization on ${\mathcal{B}}$. There is an algebra homomorphism
@@
where $T\rightarrow T$ is by taking $l$-th power, and $\widetilde{G^{*}}\rightarrow T$ is induced by $B=N\times T\rightarrow T$, $nt\mapsto t^{2}$ for any $n\in N$ and $t\in T$. Note that ${\mathcal{V}}$ is affine over ${\mathcal{B}}$. By [44, Thm 5.2], there is an isomorphism from the central subsheaf of $\widetilde{{\mathcal{D}}}_{\zeta}$ to the structure sheaf ${\mathcal{O}}_{{\mathcal{V}}}$ of ${\mathcal{V}}$. We thus also view $\widetilde{{\mathcal{D}}}_{\zeta}$ as a (coherent) sheaf of algebras on ${\mathcal{V}}$. The homomorphism ${\mathfrak{U}}_{\zeta}\rightarrow\widetilde{{\mathcal{D}}}_{\zeta}$ restricting on their centers is compatible with the natural map
@@ 3.2.2. Localization theorem
From now on, we assume moreover that $l$ is a power of prime, following [42].
@@
$R\Gamma:D^{b}(\widetilde{{\mathcal{D}}}_{\zeta}\text{-}\mathrm{mod})\rightarrow D^{b}({\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}).$
@@
$R\Gamma:D^{b}(\widetilde{{\mathcal{D}}}_{\zeta}\text{-}\mathrm{mod}_{(\chi,\zeta^{2\lambda})}){\ \xrightarrow{\sim}\ }D^{b}({\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(\chi,\chi_{\lambda})}).$

### v4

@@ 2.3.4. Translation functors
Let $R$ be a commutative Noetherian $S$-algebra. For any $\omega_{1},\omega_{2}\in\Xi$, there is a unique dominant element $\nu$ in the $W$-orbit of $\omega_{2}-\omega_{1}$. Denote by $V(\nu)$ the Weyl module for $U_{\zeta}$ of highest weight $\nu$, see e.g. [2]. We view $V(\nu)$ as a $U$-module via the natural map $U\rightarrow U_{\zeta}$. We define the translation functors by
${\mathsf{T}}_{\omega_{1}}^{\omega_{2}}:U\text{-}\mathrm{Mod}^{\Lambda,\omega_{1}}_{R}\rightarrow U\text{-}\mathrm{Mod}^{\Lambda,\omega_{2}}_{R},\quad M\mapsto\mathrm{pr}_{\omega_{2}}(M\otimes V(\nu)),$
${\mathsf{T}}^{\omega_{1}}_{\omega_{2}}:U\text{-}\mathrm{Mod}^{\Lambda,\omega_{2}}_{R}\rightarrow U\text{-}\mathrm{Mod}^{\Lambda,\omega_{1}}_{R},\quad M\mapsto\mathrm{pr}_{\omega_{1}}(M\otimes V(\nu)^{*}),$
where $\mathrm{pr}_{\omega_{i}}$ is the natural projection to the block $U\text{-}\mathrm{Mod}^{\Lambda,\omega_{i}}_{R}$. Note that ${\mathsf{T}}_{\omega_{1}}^{\omega_{2}}$ and ${\mathsf{T}}^{\omega_{1}}_{\omega_{2}}$ are exact and biadjoint to each other. The translation functors preserve the subcategory ${\mathcal{O}}^{U}_{R}$ and induce translation functors between blocks in ${\mathcal{O}}^{U}_{R}$.
For any simple reflection $s\in{\mathrm{I}}_{\mathrm{af}}$, we fix an element $\omega_{s}\in\Xi_{\mathrm{sc}}$ whose stabilizer under the $W_{\mathrm{af}}$-action is $\{1,s\}$. We abbreviate ${\mathsf{T}}^{0}_{s}={\mathsf{T}}^{0}_{\omega_{s}}$ and ${\mathsf{T}}_{0}^{s}={\mathsf{T}}_{0}^{\omega_{s}}$. Define the reflection functor on $U\text{-}\mathrm{Mod}^{\Lambda,0}_{R}$ by
$\Theta^{r}_{s}:={\mathsf{T}}^{0}_{s}\circ{\mathsf{T}}_{0}^{s}.$
We also consider the functor ${\mathbb{T}}^{r}_{s}:=\ker(\Theta^{r}_{s}\xrightarrow{\text{counit}}1)$.
@@
We can similarly define translation functors and reflection functors on the categories ${\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{\chi_{\omega}}$ with $\omega\in\Xi$.
@@ 2.3.5. $\Lambda$-translations
Let $U^{\prime}$ be $U$ as in §2.3.3, or be $u_{\zeta}$, $u_{\zeta}U^{+}_{\zeta}$. Let $R$ be a deformation ring for $U^{\prime}$. For any $\nu\in\Lambda$, there is a trivial $U^{\prime}$-module ${\mathbb{C}}_{l\nu}$ supported on the weight $l\nu$. It gives auto-equivalences $\otimes{\mathbb{C}}_{l\nu}$, $\nu\in\Lambda$ on $U^{\prime}\text{-}\mathrm{Mod}^{\Lambda}_{R}$ preserving each blocks, called the $\Lambda$-translations.
@@ 3.1. Non-commutative Springer resolutions
[…]
@@ 3.1.1. Affine braid group actions
${{\mathcal{T}}or}_{i}^{{\mathcal{O}}_{\mathfrak{g}}}({\mathcal{O}}_{\widetilde{{\mathfrak{g}}}},{\mathcal{O}}_{X})=0,\quad\forall i>0.$
We set $\widetilde{{\mathfrak{g}}}_{P,X}:=\widetilde{{\mathfrak{g}}}_{P}\times_{\mathfrak{g}}X$ and write $\pi_{P}=\pi_{P}\times_{\mathfrak{g}}X$ by abuse of notations. There is a functor $\Theta_{P}^{c}:=\pi_{P}^{*}\pi_{P*}$ on $D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})$. We abbreviate $\Theta^{c}_{s}=\Theta^{c}_{P_{s}}$ for any $s\in{\mathrm{I}}$, which are called the reflection functors.
@@
For any $s\in{\mathrm{I}}$, there is a self-equivalence ${\mathbb{T}}^{c}_{s}$ on $D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})$ with natural transformations
${\mathbb{T}}^{c}_{s}\rightarrow\Theta^{c}_{s}\xrightarrow{\text{counit}}1\xrightarrow{+1},$
which is a distinguished triangle when applying on any object in $D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})$. The assignments
$\tilde{s}\mapsto({\mathbb{T}}^{c}_{s})^{-1}\quad\text{and}\quad\theta_{\lambda}\mapsto-\otimes^{L}_{{\mathcal{O}}_{\widetilde{{\mathfrak{g}}}_{X}}}{\mathcal{O}}_{\widetilde{{\mathfrak{g}}}_{X}}(\lambda),\qquad\forall s\in{\mathrm{I}},\ \forall\lambda\in\Lambda,$
define a (weak, right) $\mathfrak{B}_{\mathrm{ex}}$-action on $D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})$. Moreover, if $X\rightarrow{\mathfrak{g}}$ is equivariant for a closed subgroup $K\subset G\times{\mathbb{C}}^{\times}$, the constructions above are naturally $K$-equivariant.
@@ 3.1.2. Non-commutative Springer resolutions
Let $\mathfrak{B}^{+}_{\mathrm{ex}}$ be the semi-group of $\mathfrak{B}_{\mathrm{ex}}$ generated by the canonical lifts $\tilde{s}$ with $s\in{\mathrm{I}}_{\mathrm{af}}$.
@@
    There is a unique t-structure (called the exotic t-structure) on $D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})$ defined by
    $D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})^{\leq}=\{{\mathcal{F}}|\ \pi_{*}(b{\mathcal{F}})\in D^{\leq}\mathrm{Coh}X,\ \forall b\in\mathfrak{B}^{+}_{\mathrm{ex}}\}$;
    $D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})^{\geq}=\{{\mathcal{F}}|\ \pi_{*}(b^{-1}{\mathcal{F}})\in D^{\geq}\mathrm{Coh}X,\ \forall b\in\mathfrak{B}^{+}_{\mathrm{ex}}\}$.
@@
    There exists a $G\times{\mathbb{C}}^{\times}$-equivariant tilting bundle $\mathcal{E}$} on $\widetilde{{\mathfrak{g}}}$ such that
    $R\mathcal{H}om_{\widetilde{{\mathfrak{g}}}_{X}}({\mathcal{E}}_{X},-):\ D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})\rightarrow D^{\mathrm{b}}({\mathsf{A}}_{X}\text{-}\mathrm{mod})$
    is an equivalence, where ${\mathcal{E}}_{X}:={\mathcal{E}}\times_{\mathfrak{g}}X$, ${\mathsf{A}}_{X}={\mathsf{A}}\otimes_{{\mathcal{O}}_{\mathfrak{g}}}{\mathcal{O}}_{X}$ and ${\mathsf{A}}=\mathcal{E}nd_{\widetilde{{\mathfrak{g}}}}^{\mathrm{op}}({\mathcal{E}})$, and such that this equivalence is t-exact with respect to the exotic t-structure on the LHS and the trivial t-structure on the RHS.
@@
    Moreover, if $X\rightarrow{\mathfrak{g}}$ is equivariant for a closed subgroup $K\subset G\times{\mathbb{C}}^{\times}$, we have similar statements for $D^{\mathrm{b}}\mathrm{Coh}^{K}(\widetilde{{\mathfrak{g}}}_{X})$.
By definition, the functor $\pi_{*}$ is t-exact for the exotic t-structure on $D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})$. It yields a functor $\pi_{*}:{\mathsf{A}}_{X}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}(X)$. By [11], the functors $\Theta^{c}_{P}$, $\Theta^{c}_{s}$ ($s\in{\mathrm{I}}_{\mathrm{af}}$) are also exact for the exotic t-structure. We thus view them as functors on ${\mathsf{A}}_{X}\text{-}\mathrm{mod}$.
@@
[…]
@@
[…]
@@ 3.2. Localization theorem for ${\mathfrak{U}}_{\zeta}$
The localization theorem for ${\mathfrak{U}}_{\zeta}$ was established by Backelin–Kremnizer [4] and Tanisaki [42, 43, 39]. In this subsection, we give a reminder on the localization theorem mainly following [39].
@@ 3.2.1. Quantum differential operators
[…]
@@
where $T\rightarrow T$ is by taking $l$-th power, and $\widetilde{G^{*}}\rightarrow T$ is induced by $B=N\times T\rightarrow T$, $nt\mapsto t^{2}$ for any $n\in N$ and $t\in T$. Note that $\mathcal{V}$} is affine over $\mathcal{B}$}. By [42, Thm 5.2], there is an isomorphism from the central subsheaf of $\widetilde{{\mathcal{D}}}_{\zeta}$ to the structure sheaf ${\mathcal{O}}_{{\mathcal{V}}}$ of $\mathcal{V}$}. We thus also view $\widetilde{{\mathcal{D}}}_{\zeta}$ as a (coherent) sheaf of algebras on $\mathcal{V}$}. The homomorphism ${\mathfrak{U}}_{\zeta}\rightarrow\widetilde{{\mathcal{D}}}_{\zeta}$ restricting on their centers is compatible with the natural map
@@ 3.2.2. Localization theorem
From now on, we assume moreover that $l$ is a prime power, following [39].
@@
$R\Gamma:D^{\mathrm{b}}(\widetilde{{\mathcal{D}}}_{\zeta}\text{-}\mathrm{mod})\rightarrow D^{\mathrm{b}}({\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}).$
@@
$R\Gamma:D^{\mathrm{b}}(\widetilde{{\mathcal{D}}}_{\zeta}\text{-}\mathrm{mod}_{(\chi,\zeta^{2\lambda})}){\ \xrightarrow{\sim}\ }D^{\mathrm{b}}({\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(\chi,\chi_{\lambda})}).$